OpenStudy (anonymous):

Integral 1+root(u)/root(u)

OpenStudy (zzr0ck3r):

\[\int\frac{1+\sqrt{u}}{\sqrt{u}}du\]?

OpenStudy (anonymous):

\[\int\limits_{9}^{4}\frac{ 1+\sqrt{u} }{ \sqrt{u} }\]

OpenStudy (zzr0ck3r):

\[\int_{9}^{4}\frac{1+\sqrt{u}}{\sqrt{u}}du=\int_{9}^{4}u^{-\frac{1}{2}}du+\int_{9}^{4}du\] make sense?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

can you do it now?

OpenStudy (anonymous):

so just integrate each?

OpenStudy (zzr0ck3r):

yep